monosodium glutamate (MSG), a flavor enhancer in certain foods containing 35.51% carbon (atomic mass=12.01 g/mol) 4.77% hydrogen (atomic mass=1.01 g/mol) 37.85% oxygen (ato,ic mass=16.00 g/mol) 8.29 nitrogen (atomic mass= 14.01 g/mol) and 13.60% sodium (atomic mass= 22.99 g/mol). calculate the chemical formula of MSG.


1
Expert's answer
2021-11-03T08:13:48-0400

For example, 100 g of MSG will contain 35.51/12.01 = 2.956 mol of C, 4.77/1.007 = 4.726 mol of H, 37.85/15.99 = 2.367 mol of O, 8.29/14.006 = 0.5918 mol of N, 13.60/22.98 = 0.5918 mol of Na.


Then all quantities will be divided by the smallest amount of moles.



So, the empirical formula is C5H8O4NNa.


The molar mass of the empirical formula will be calculated.

"(5\u00d712.01)+(8\u00d71.007)+(4\u00d715.99)+14.006+22.98=169.05g\/mol"

The molecular formula of MSG is the same as the empirical formula C5H8O4NNa.